OpenStudy (anonymous):

solve for k: ln(k+1)-ln(k)=.125

OpenStudy (anonymous):

Properties of logs: ln(x)-ln(y)=ln(x/y) so: ln(k+1)-ln(k)=ln((k+1)/k)=ln(1+1/k) Remember e^ln(x)=x. So if you make all terms e^(something) it makes it much easier: => e^ln(1+1/k)=e^0.125 => 1+1/k=e^0.125 => 1/k=e^0.125-1 => k=1/(e^0.125-1) => k= 7.51041...

OpenStudy (anonymous):

thanks so much